A wire is stretched to n times it's length. What happens to its resistance?
1
Expert's answer
2021-07-12T12:17:41-0400

First of all, the resistance is proportional to the quantity "\frac{l}{S}", where "S" is a cross-section area and "l" is wire's length. Secondly, let us remark that the wire's total volume, "V=S\cdot l", in conserved during the stretching. Therefore, if we have stretch its length by "n" times, we have reduced its cross-section area by "n" times. From this we conclude that the resistance is increased by "\frac{n}{1/n}=n^2" times.